OpenStudy (anonymous):

Hi! I'm new to open study can anyone Please help me with this problem :) 1) find the equation of the line that passes through (-3,-3) and is parallel to x-2y=-7

OpenStudy (amoodarya):

welcome to openstudy \[y-y_0=m(x-x_0)\] a line passes (x0,y0) with slope m

OpenStudy (anonymous):

thank you so much :)

OpenStudy (amoodarya):

\[(x_0,y_0)=(-3,-3)\\x-2y=-7 \rightarrow -2y=-x-7 \\y=\frac{-x-7}{-2}\\y=\frac{1}{2}x+\frac{7}{2} \rightarrow m=\frac{1}{2}\] can you go on ? let me see your work

OpenStudy (anonymous):

Im confused on how you got 1/2x :( Can you guide me through so I can understand it better? Math was never my strong points and I have no clue what Im doing

OpenStudy (amoodarya):

\[y=\frac{ -x-7 }{ -2 }=\frac{ -(x+7) }{ -(2) }=\frac{ x+7 }{2 }=\\\frac{ x }{2 }+\frac{ 7 }{2 }\]

OpenStudy (anonymous):

ohhh ok that makes more sense now :)

OpenStudy (amoodarya):

can you go on ?